Section 14  Mathematical Aspects of Computer Science

Section 14: Mathematical Aspects of Computer Science

Invited Lectures

Wednesday, 19 August  H 2032

14:00 – 14:45 Joan Feigenbaum, AT&T Labs Research, Florham Park, USA Games, Complexity Classes, and Approximation Algorithms (I: p.167) Chair: Hans J. Prömel

15:00 – 15:45 Toniann Pitassi, University of Arizona, Tucson, USA Combinatorial and algebraic techniques in proof complexity (I: p.171) Chair: Helmut Alt

Thursday, 20 August  H 2032

16:15 – 17:00 Madhu Sudan, MIT, Cambridge, USA Probabilistic checking of proofs (I: p.173) Chair: Hans J. Prömel

17:15 – 18:00 Johan Håstad, Royal Institute of Technology, Stockholm, Sweden On approximating NP-hard optimization problems (I: p.169) Chair: Hans J. Prömel

Friday, 21 August  H 2032

14:00 – 14:45 Emo Welzl, ETH Zürich, Zürich, Switzerland Halving Point Sets (I: p.175) Chair: Helmut Alt

15:00 – 15:45 Miklós Ajtai, IBM Almaden Research Center, San Jose, USA Worst-case Complexity, Average-case Complexity, and Lattice Problems (I: p.165) Chair: Helmut Alt

Short Communications and Poster Sessions

Short Communications  Thursday, 20 August
Mathematical Aspects of Computer Science I  MA 043

Chair: Stefan Felsner

14:00 – 14:15 Laxminarayan Das, Texas A&M University, College Station, Texas, USA Convexity to Invexity (p.290)

14:15 – 14:30 Narayan C. Debnath, Winona State University, Winona, MN, USA RGPG: A Graph for Software Modeling (p.290)

14:30 – 14:45 Victor Kasyanov, Institute of Informatics Systems, Novosibirsk, Russia Hierarchical graphs and visual processing (p.292)

Mathematical Aspects of Computer Science  Section 14

Mathematical Aspects of Computer Science II  MA 043

Chair: Clemens Gröpl

15:00 – 15:15 Guy Merlin Mbakop, Institut für Mathematik, Humboldt–Universität, Berlin, Germany An Efficient Algorithm for Solving Real Polynomial Equations: The Complete Intersection Case

(p.294)

15:15 – 15:30 Bernard Mourrain, INRIA, SAGA, Sophia Antipolis, France Multivariate polynomial equations and duality (p.294)

15:30 – 15:45 Emil J Volcheck, Baltimore, Maryland, USA Computing with algebraic curves: a survey of recent results (p.295)

Short Communications  Friday, 21 August
Mathematical Aspects of Computer Science III  MA 043

Chair: Clemens Gröpl

16:15 – 16:30 Barbu Gheorghe, University of Piteşti, Pitesti, Romania A method for Computer Generation of Random Vectors (p.289)

16:30 – 16:45 Sven Hartmann, Universität Rostock, Rostock, Germany New results on the consistency of cardinality constraints (p.292)

16:45 – 17:00 Stefan Felsner, Freie Universität Berlin, Berlin, Germany The Linear-Extension-Diameter (p.291)

Mathematical Aspects of Computer Science IV  MA 043

Chair: Stefan Felsner

17:15 – 17:30 Svetlana Asmuss, University of Latvia, Riga, Latvia On optimal algorithms of approximation under imprecise information (p.289)

17:30 – 17:45 Clemens Gröpl, Humboldt–Universität zu Berlin, Berlin, Germany Random OBDDs (p.292)

17:45 – 18:00 Andrew Adamatzky, University of the West of England, Bristol, United Kingdom On computation universality of excitable lattices

Short Communications  Saturday, 22 August
Mathematical Aspects of Computer Science V  MA 043

Chair: Hartmut Ehrig

14:00 – 14:15 Bernd Gärtner, ETH Zürich, Zürich, Switzerland Abstract Objective Functions — From Geometry to Combinatorics in Linear Programming (p.291)

14:15 – 14:30 Andreas Brieden, Technische Universität München, Munich, Germany Optimal Radii-Algorithms (p.289)

14:30 – 14:45 Anuar Dusembaev, National University of Kasakhstan, Almaty, Kasachstan An operator approach to program segmentation (p.290)

Mathematical Aspects of Computer Science VI  MA 043

Chair: Andreas Brieden

15:00 – 15:15 Hartmut Ehrig, Technische Universität Berlin, Berlin, Germany On the Role of Category Theory in Computer Science (p.290)

15:15 – 15:30 Raja Natarajan, Tata Institute of Fundamental Research, Bombay, India Combinators for Concurrency (p.294)

15:30 – 15:45 George Francis, University of Illinois, Urbana, IL, USA The Making of the Optiverse (p.291)

2  ICM’98 – Program

Section 14  Mathematical Aspects of Computer Science

Poster Session  Saturday, 22 August 16:15 – 18:00  MA 144

Giovanni Francesco Mascari, Coniglio Nazionale delle Ricerche, Rome, Italy

From TQFT to TQCT (p.293) Roman Matuszewski, University of Białystok, Białystok, Poland

Journal of Formalized Mathematics — a collection of mechanically checked articles (p.293)

ICM’98 – Program  3